If there are twelve small squares, four small squares are shaded, how many more small squares need to be shaded so that 2/3 of the rectangle is shaded?

a. 2

b. 3

c. 4

d. 5


1
Expert's answer
2021-05-10T17:15:14-0400

If there are 12 small squares, and 4 small squares are shaded, then to shade 2/3 = 8/12 of the rectangle we need 8 - 4 = 4 more small squares.

So, the correct answer is c. 4.